The fifth term in an arithmetic sequence is 110 and the seventeenth term is 74. Find the tenth term in the sequence.



1
Expert's answer
2021-02-03T15:24:30-0500

Given an arithmetic sequence


"a_5=a+4d=110"

"a_{17}=a+16d=74"

Then


"a+16d-(a+4d)=74-110"

"12d=-36"

"d=-3"

Find "a"

"a+4(-3)=110"

"a=122"

Hence


"a_{10}=a+9d=122+9(-3)=95"

The tenth term in the sequence is "95."